the uncertainty associated with a temperature calculated from two readings of a 0oC and 110oC thermometer is 1oC.calculate the percentage uncertainty associated with a temperature difference of 8.0oC. give your answer to 2 significant figures


1
Expert's answer
2022-03-28T19:18:02-0400

The uncertainty between 0o C and 110o C is 1o C

"\frac{\delta ^oC}{T}= \frac{1}{110}= 0.0090"

for 8 oC temperature difference = "8\times 0.0090= 0.0727 ^oC"